A part of a storeroom measures 14 1/2 feet by 6 1/4 feet. The number of square feet in this part is?
Answer by KnightOwlTutor(293)

You would multiply the two values together
141/2=29/2
61/4=25/4
(29/2)(25/4=725/8=90.625 square feet